OpenStudy (anonymous):

can someone please explain to me how the derivative of f(x)=sin x is?? and how its found.

OpenStudy (amistre64):

lim as h-> 0; sin(x+h)-sin(x) ------------ h

OpenStudy (anonymous):

so, theres no cosine involved in the answer.

OpenStudy (amistre64):

sin(x+h) = sin(x)cos(h) + sin(h)cos(x) sin(x)cos(h) + sin(h)cos(x) - sin(x) = sin(x)(cos(h)-1) + sin(h)cos(x)

OpenStudy (amistre64):

\[sin(x)\frac{cos(h)-1}{h}+cos(x)\frac{sin(h)}{h}\]

OpenStudy (amistre64):

\[\frac{cos(h)-1}{h} \implies 0;\ and \frac{sin(h)}{h} \implies 1\]

OpenStudy (amistre64):

0 + cos(x) = cos(x) Dx(sin(x)) = cos(x)

OpenStudy (anonymous):

\[d(\sin x)/dx=\lim_{dx \rightarrow 0} [\sin(x+dx)- \sin(x)]/[x+dx-x]\] \[d(\sin x)/dx=\lim_{dx \rightarrow 0} [2\cos(x+dx/2)*\sin(dx/2)]/[dx]\] \[= \lim_{dx \rightarrow 0} [\cos(x+dx/2)]*\ [\sin(dx/2)]/[dx/2]\] =cos(x)

OpenStudy (anonymous):

in the change of delta x, why is h by itself , what were the other two variables that were cancelled out

OpenStudy (anonymous):

go for it bro dont be shamed

OpenStudy (amistre64):

x+h-x .... = h

OpenStudy (anonymous):

\[\lim_{a \rightarrow 0} \sin(a)/a=1\]

OpenStudy (anonymous):

i just cant understand it how sin x ends up cos x , the derivative of the function,

OpenStudy (anonymous):

sin(A)-sin(B)=2*cos((A+B)/2)*sin((A-B)/2)

OpenStudy (anonymous):

hope you know the general eqn for finding derivatives...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

no , LOL, ill get it though thats a promise